OpenStudy (anonymous):

solve the equation -2y+9=11

OpenStudy (bahrom7893):

-2y=2 y=-1

OpenStudy (anonymous):

-2y + 9 = 11 - 9 -9 -2y = 2 /-2 /-2 y = -1

OpenStudy (anonymous):

2y =9 - 11 2y = -2 y = -1